# Surface Tension of a Topological Phase

### v2

@@ I.2 Contribution of bulk and edge states (Numerical)
$\Delta E=\Delta E_{\text{bulk}}+\Delta E_{\text{edge}},$ (S10)
@@
$\Delta E_{\rm{bulk}}=\sum_{j=1}^{L^{2}-l}\Delta E_{j}=4\Delta\sigma_{\rm{bulk}}\sqrt{n},$ (S11a)
$\Delta E_{\text{edge}}=\sum_{j=L^{2}-l+1}^{L^{2}}\Delta E_{j}=4\Delta\sigma_{\text{edge}}\sqrt{n}.$ (S11b)
@@
Thus as a function of $\kappa$, i.e. the coupling between the spin and the fermion, the complete surface tension $\Delta\sigma$ can be separately decomposed into its bulk and edge contributions where
$\Delta\sigma=\Delta\sigma_{\text{bulk}}+\Delta\sigma_{\text{edge}}.$ (S12)
@@
While in the main text, the variation of $\Delta\sigma$ was shown as a function of $\kappa$, in Fig. S2(a) we show the individual contributions from both the bulk and edge states. Deep in the topological regime $\kappa\sim 1$, one finds that the contribution to $\Delta\sigma$ is dominantly from the edge states. This is reflected even in the $\Delta E_{j}$, where we show its variation as a function of the eigenvalue index $j$ for $\kappa=1.2$, which shows that only $\sim l$ states near the Fermi energy have significant $\Delta E_{j}$ (see Fig. S2(b)). In contrast, when $\kappa=2.5$ such that the system is in a trivial phase (no edge states on the droplet boundary), the sole contribution to $\Delta\sigma$ is due to the bulk states. The behavior of $\Delta E_{j}$ is non-universal and fluctuates over all the bulk states in the trivial phase (see Fig. S2(c)).
@@
A comparison of the $\Delta\sigma_{\text{edge}}$ with the analytic estimate $\Delta{\mathcal{E}}_{\rm{edge}}$ from the last section (where we set $\hbar=1$) is shown in Fig. S2(d), illustrating the topological origin of the surface tension in the regime $20<|\kappa|<$ 2 which is universal and applicable to any topological system with a boundary manifold. Any contribution to $\Delta\sigma$ in the trivial phase $|\kappa|>2$ is due to the bulk states, which is model dependent and non-universal as discussed in the section below.
II Non-Universal Surface Tension of Trivial Insulators
@@
In this section, we discuss how the contribution to $\Delta\sigma_{\text{bulk}}$ is non-universal and model dependent using two different Hamiltonian systems.
II.1 Trivial phases in BHZ model
@@
Figure S3: Surface tension for various filling fractions in trivial situation: (a) Difference of energy $\Delta E_{j}$ of $j$-th single-particle level (see Eq. (S8)) after and before the quench in trivial situation ($\kappa=4.5$) when there is a square shaped droplet of size $n=l^{2}$ (where $l=16$) with $s_{i}=1$, while the rest of $s_{i}=-1$. (b) Surface tension $\Delta\sigma_{\rm{bulk}}^{f}$ as a function of filling fraction $f$ for various values of $\kappa$ in trivial situation of BHZ model. (c) $\Delta E_{j}$ as a function of $j$ in trivial insulator with $\kappa^{\prime}=4.5$ (see Hamiltonian in Eq. (S14)) when there is a square shaped droplet of size $n=l^{2}$ (where $l=16$). (d) $\Delta\sigma_{\rm{bulk}}^{f}$ as a function of filling fraction $f$ for various values of $\kappa^{\prime}$ in trivial insulator . In all plots, number of unit cells is $L^{2}=4096$.
@@
We first analyze the variation of the surface tension in trivial situation ($|\kappa|>2$) of the model discussed in the main text, where the $\Delta E_{j}$ for $\kappa=2.5$ was shown in Fig. S2(c). In Fig. S3(a) we show $\Delta E_{j}$ for $\kappa=4.5$, again reflecting the non-universal fluctuation of $\Delta E_{j}$ about zero. Furthermore, we find that both the value and the sign of $\Delta\sigma_{\text{bulk}}$ can be filling ($f$) dependent. In Fig. S3(b), we plot $\Delta\sigma_{\text{bulk}}^{f}$ using
$\Delta E(f)=\sum_{j=1}^{j_{\rm{max}}}\Delta E_{j}=4(\Delta\sigma^{f}_{\text{bulk}})\sqrt{n},$ (S13)
@@
II.2 Quadratic trivial band insulator
@@
$H=-\kappa^{\prime}\sum_{i}s_{i}{\bf{\Psi}}_{i}^{\dagger}\sigma_{z}~{}{\bf{\Psi}}_{i}-\frac{1}{2}\sum_{\langle ij\rangle}({\bf{\Psi}}_{i}^{\dagger}\sigma_{z}~{}{\bf{\Psi}}_{j}+{\rm{h.c.}})-J\sum_{\langle ij\rangle}s_{i}s_{j}-h\sum_{i}s_{i},$ (S14)
@@
where ±$\pm\kappa^{\prime}s_{i}$ is the staggered mass of $A$ and $B$ sites of $i$-th unit cell and the hopping is restricted only between $A$-$A$ and $B$-$B$ sites of the nearest neighboring unit cells. For uniform $s_{i}=1$, we assume $M^{\prime}\equiv\kappa^{\prime}$ and the Hamiltonian can be recast in momentum space for $k_{x},k_{y}\in[-\pi,\pi)$ as
$H_{k_{x},k_{y}}=(-M^{\prime}-\cos(k_{x})-\cos(k_{y}))\sigma_{z}.$ (S15)
@@
Thus, the bulk energy-gap is zero for $|M^{\prime}|\leq 2$ where the system behaves as a metal, while for $|M^{\prime}|>2$ the bulk energy-gap is finite and the system is a trivial insulator. Just like the model presented in the main text the spectrum has a $M^{\prime}\to-M^{\prime}$ symmetry but at $M^{\prime}=\pm 2$ the metallic phase gives way to an insulator at half-filling via quadratically dispersing band. Importantly, there is no topological phase in this model. We now consider the same quench protocol as performed for BHZ model discussed in the main text. For a trivial insulator ($|\kappa^{\prime}|\gg 2$), we find that $\Delta E_{j}$ for all the bulk states shows non-universal fluctuation about zero (see Fig. S3(c)). Similar to the trivial phase of BHZ model, here also $\Delta\sigma_{\rm{bulk}}^{f}$ shows non-universal dependence on filling $f$ and $\Delta\sigma_{\rm{bulk}}^{f}<0$ at half-filling (see Fig. S3(d)).
@@
Therefore, the surface tension $\Delta\sigma_{\rm{bulk}}$ in trivial phase at half-filling (i.e. ground state) can be either positive or negative which is decided by the random contributions from individual single-particle energy-levels, rendering the behavior of surface tension in trivial situation non-universal.
III Role of Symmetries and Topological Surface Tension
In this section, we demonstrate the role of symmetries in affecting the surface tension in topological phases. The model discussed in the main text has a set of intricate symmetries such that the contribution of the surface tension can be distilled. These are enumerated below:
  1. 1.
    Spin-fermion flip symmetry: The energy eigenvalues $E_{j}$ of bulk spectrum are independent of sign of $\kappa$, i.e. $E_{j}(\kappa)=E_{j}(-\kappa)$. This also reflects into a spin-flip symmetry as $E_{j}(\kappa,\{s_{i}\})=E_{j}(\kappa,\{-s_{i}\})$. This guarantees that under the quench protocol, the bulk energy density contribution $\Delta\epsilon$ to free energy equals zero. Microscopically this corresponds to a symmetry operation $s_{i}\rightarrow-s_{i},{\bf{\Psi}}_{i}\rightarrow(-1)^{i_{x}+i_{y}}\sigma_{x}{\bf{\Psi}}_{i}\otimes K$ where $K$ is the complex conjugation and $(i_{x},i_{y})$ is the position of $i$-th unit cell.
  2. 2.
Note that the fermion Hamiltonian does not have microscopic time-reversal symmetry - which results in stabilization of finite Chern number phases. In the following, we examine the role of breaking each of these symmetries.
@@ III Role of Symmetries and Topological Surface Tension
III.1 Role of breaking spin-fermion flip symmetry (class D model)
@@
Figure S4: Effect of hopping between next nearest neighbors: (a) Chern number $C$ as a function of mass parameter $M$, (b) change in energy density $\Delta\epsilon$ as a function of $\kappa$, (c) additional surface tension $\Delta\sigma$ as a function of $\kappa$, (d) enhancement ratio $\gamma=n_{c}^{\rm{Q}}/n_{c}^{\rm{cl}}$ of critical cluster size as a function of $\kappa$ when the hopping between next nearest neighboring unit cells is absent ($t_{2}=0.0$) and present ($t_{2}=1.0$). In all plots, $J=0.2$, $h=1.0$, $T=0$, $L^{2}=4096$.
@@
In the previous section we had pointed out that the Hamiltonian in the main text has a symmetry where the sign of the spin does not change the single particle energy eigenvalues. In order to break this symmetry (while retaining particle-hole symmetry), we introduce the next-nearest neighbor hopping $t_{2}$ such that the complete spin-fermion hopping Hamiltonian becomes
$H=-\kappa\sum_{i}s_{i}{\bf{\Psi}}_{i}^{\dagger}\sigma_{z}~{}{\bf{\Psi}}_{i}-\sum_{\langle ij\rangle}({\bf{\Psi}}_{i}^{\dagger}{\large{\eta}}_{ij}~{}{\bf{\Psi}}_{j}+{\rm{h.c.}})-t_{2}\sum_{\langle\langle ij\rangle\rangle}({\bf{\Psi}}_{i}^{\dagger}{\large{\eta}}^{\prime\prime}_{ij}~{}{\bf{\Psi}}_{j}+{\rm{h.c.}})-J\sum_{\langle ij\rangle}s_{i}s_{j}-h\sum_{i}s_{i},$ (S16)
@@
where $t_{2}$ is the strength of the hopping between the next nearest neighboring unit cells and $\eta^{\prime\prime}_{ij}=\frac{1}{2}(\sigma_{z}+i\cos(\theta_{ij})\sigma_{x}+i\sin(\theta_{ij})\sigma_{y})$, the angle $\theta_{ij}$ between the $x$-axis and the bond connecting the next nearest neighboring unit cells can assume the values $\theta_{ij}=\pi/4$ and $\theta_{ij}=3\pi/4$. For uniform $s_{i}$ throughout the system, we assume $M\equiv\kappa s_{i}$ and the Hamiltonian in momentum space for $k_{x},k_{y}\in[-\pi,\pi)$ is written as
@@
$\sigma_{x}H^{*}_{k_{x},k_{y}}\sigma_{x}=-H_{-k_{x},-k_{y}}.$ (S18)
For uniform $s_{i}$, the system can host topological phases having Chern numbers $C=\pm 2,\pm 1$ depending on the value of mass parameter $M$ when $t_{2}\neq 0$ (see Fig. S4(a)). Interestingly, the ground state energy of the fermionic system, which is given by
[…]
where $\text{where~{}}E_{k_{x},k_{y}}$ $=-\sqrt{d_{1}^{2}+d_{2}^{2}+d_{3}^{2}}~{},$ (S19)
is not symmetric under $M\to-M$, i.e., $E_{g}(M)\neq E_{g}(-M)$ when $t_{2}\neq 0$. Thus, the sudden quench from $h<0$ to $h>0$ causes an additional correction to bulk energy density $\Delta\epsilon$ along with the correction to surface tension $\Delta\sigma$ when $t_{2}\neq 0$, such that the change in the energy in the fermionic system for a formation of a square-shaped droplet of size $n$ is
$\Delta E=(\Delta\epsilon)n+4\sqrt{n}(\Delta\sigma).$ (S20)
We show the variation of $\Delta\epsilon$ and $\Delta\sigma$ in Fig. S4(b,c), where we observe that unlike the situation with $t_{2}=0$, $\Delta\epsilon\neq 0$ for $\kappa\neq 0$ when $t_{2}\neq 0$. Furthermore, in the topological situation $20<|\kappa|<$ 2, surface tension $\Delta\sigma$ for $t_{2}\neq 0$ is greater than that for $t_{2}=0$. This can be understood as follows: when $t_{2}\neq 0$ and $20<\kappa<$ 2, a droplet of $s_{i}=1$ corresponds to Chern number $C=2$, while its background with $s_{i}=-1$ has $C=-1$. Thus, the difference in Chern numbers ($|\Delta C|=3$) between the droplet and its background for $t_{2}\neq 0$ is greater than $|\Delta C|=2$ for $t_{2}=0$. An increased $|\Delta C|$ in topological situation causes greater $\Delta\sigma$ since the number of edge modes increases, increasing the effective $\Delta{\mathcal{E}}_{\text{edge}}$. This results in a higher surface tension for $t_{2}\neq 0$ compared to that for $t_{2}=0$. For $-2<\kappa<0$, $\Delta\sigma$ shows similar behavior where a droplet of $s_{i}=1$ has $C=-1$ and its background with $s_{i}=-1$ has $C=2$. In presence of non-zero $\Delta\epsilon$ and $\Delta\sigma$, the critical cluster size for $t_{2}\neq 0$ modifies to
$n_{c}^{\rm{Q}}=\frac{\pi(\sigma_{\rm{cl}}+\Delta\sigma)^{2}}{(2|h|-\Delta\epsilon)^{2}},$ (S21)
@@
$\gamma=\frac{n_{c}^{\rm{Q}}}{n_{c}^{\rm{cl}}}=\frac{(1+(\Delta\sigma/\sigma_{\rm{cl}}))^{2}}{(1-(\Delta\epsilon/(2|h|)))^{2}}.$ (S22)
We find that in topological situation, $\gamma$ can be greater than $1$ for a suitable choice of $J$ and $h$, indicating an enhancement of critical cluster size (see Fig. S4(d)).
III.2 Role of breaking particle-hole symmetry (class A model)
Figure S5: Effect of particle-hole symmetry breaking hopping term: (a) Chern number $C$ as a function of mass parameter $M$, (b) change in energy density $\Delta\epsilon$ as a function of $\kappa$, (c) additional surface tension $\Delta\sigma$ as a function of $\kappa$, (d) enhancement ratio $\gamma=n_{c}^{\rm{Q}}/n_{c}^{\rm{cl}}$ of critical cluster size as a function of $\kappa$ when particle-hole symmetry breaking hopping term is absent ($\lambda=0.0$) and present ($\lambda=0.6$). Dotted lines denote the critical points $\kappa=0,\pm 2$ when $\lambda=0$ and $\kappa=0,\pm 1.6$ when $\lambda=0.6$. In all plots, $J=0.2$, $T=0$, $L^{2}=4096$.

### v3

@@ A.2 Contribution of bulk and edge states to interface energy
$\Delta E=\Delta E_{\text{bulk}}+\Delta E_{\text{edge}},$ (23)
@@
$\Delta E_{\rm{bulk}}=\sum_{j=1}^{L^{2}-l}\Delta E_{j}=4\Delta\sigma_{\rm{bulk}}\sqrt{n},$ (24a)
$\Delta E_{\text{edge}}=\sum_{j=L^{2}-l+1}^{L^{2}}\Delta E_{j}=4\Delta\sigma_{\text{edge}}\sqrt{n}.$ (24b)
@@
Thus as a function of $\kappa$, i.e. the coupling between the spin and the fermion, $\Delta\sigma$ can be separately decomposed into its bulk and edge contributions where
$\Delta\sigma=\Delta\sigma_{\text{bulk}}+\Delta\sigma_{\text{edge}}.$ (25)
@@
While in the main text, the variation of $\Delta\sigma$ was shown as a function of $\kappa$, in Fig. 8(a) we show the individual contributions from both the bulk and edge states. Deep in the topological regime $\kappa\sim 1$, one finds that the contribution to $\Delta\sigma$ is dominantly from the edge states. This is reflected even in the $\Delta E_{j}$, where we show its variation as a function of the eigenvalue index $j$ for $\kappa=1.2$, which shows that only $\sim l$ states near the Fermi energy have significant $\Delta E_{j}$ (see Fig. 8(b)). In contrast, when $\kappa=2.5$ such that the system is in a trivial phase (no edge states on the droplet boundary), the sole contribution to $\Delta\sigma$ is due to the bulk states. The behavior of $\Delta E_{j}$ is non-universal and fluctuates over all the bulk states in the trivial phase (see Fig. 8(c)).
@@
A comparison of the $\Delta\sigma_{\text{edge}}$ with the analytic estimate $\Delta{\mathcal{E}}_{\rm{edge}}$ from the last section (where we set $\hbar=1$) is shown in Fig. 8(d), illustrating the topological origin of the interface energy in the regime $0<|\kappa|<2$ which is universal and applicable to any topological system with a boundary manifold. Any contribution to $\Delta\sigma$ in the trivial phase $|\kappa|>2$ is due to the bulk states, which is model dependent and non-universal as discussed in the section below.
A.3 Non-Universal Interface Energy Contribution in Trivial Insulators
@@
We now discuss how the contribution to $\Delta\sigma_{\text{bulk}}$ is non-universal and model dependent using two different Hamiltonian systems.
@@
Figure 9: Correction to interfacial line energy for various filling fractions in trivial situation: (a) Difference of energy $\Delta E_{j}$ of $j$-th single-particle level (see Eq. (21)) after and before the quench in trivial situation ($\kappa=4.5$) when there is a square shaped droplet of size $n=l^{2}$ (where $l=16$) with $s_{i}=1$, while the rest of $s_{i}=-1$. (b) Correction to interfacial line energy $\Delta\sigma_{\rm{bulk}}^{f}$ as a function of filling fraction $f$ for various values of $\kappa$ in trivial situation of BHZ model. (c) $\Delta E_{j}$ as a function of $j$ in trivial insulator with $\kappa^{\prime}=4.5$ (see Hamiltonian in Eq. (A.3.2)) when there is a square shaped droplet of size $n=l^{2}$ (where $l=16$). (d) $\Delta\sigma_{\rm{bulk}}^{f}$ as a function of filling fraction $f$ for various values of $\kappa^{\prime}$ in trivial insulator . In all plots, number of unit cells is $L^{2}=4096$.
A.3.1 Trivial phases in BHZ model
@@
We first analyze the variation of the interfacial line energy correction in trivial situation ($|\kappa|>2$) of the model discussed in the main text, where the $\Delta E_{j}$ for $\kappa=2.5$ was shown in Fig. 8(c). In Fig. 9(a) we show $\Delta E_{j}$ for $\kappa=4.5$, again reflecting the non-universal fluctuation of $\Delta E_{j}$ about zero. Furthermore, we find that both the value and the sign of $\Delta\sigma_{\text{bulk}}$ can be filling ($f$) dependent. In Fig. 9(b), we plot $\Delta\sigma_{\text{bulk}}^{f}$ using
$\Delta E(f)=\sum_{j=1}^{j_{\rm{max}}}\Delta E_{j}=4(\Delta\sigma^{f}_{\text{bulk}})\sqrt{n},$ (26)
@@
A.3.2 Quadratic trivial band insulator
@@
$H=-\kappa^{\prime}\sum_{i}s_{i}{\bf{\Psi}}_{i}^{\dagger}\sigma_{z}~{\bf{\Psi}}_{i}-\frac{1}{2}\sum_{\langle ij\rangle}({\bf{\Psi}}_{i}^{\dagger}\sigma_{z}~{\bf{\Psi}}_{j}+{\rm{h.c.}})$
$-J\sum_{\langle ij\rangle}s_{i}s_{j}-h\sum_{i}s_{i},$ (27)
@@
where $\pm\kappa^{\prime}s_{i}$ is the staggered mass of $A$ and $B$ sites of $i$-th unit cell and the hopping is restricted only between AA-AA and $B$-$B$ sites of the nearest neighboring unit cells. For uniform $s_{i}=1$, we assume $M^{\prime}\equiv\kappa^{\prime}$ and the Hamiltonian can be recast in momentum space for $k_{x},k_{y}\in(-\pi,\pi]$ as
$H_{k_{x},k_{y}}=(-M^{\prime}-\cos(k_{x})-\cos(k_{y}))\sigma_{z}.$ (28)
@@
Thus, the bulk energy-gap is zero for $|M^{\prime}|\leq 2$ where the system behaves as a metal, while for $|M^{\prime}|>2$ the bulk energy-gap is finite and the system is a trivial insulator. Just like the model presented in the main text the spectrum has a $M^{\prime}\to-M^{\prime}$ symmetry but at $M^{\prime}=\pm 2$ the metallic phase gives way to an insulator at half-filling via quadratically dispersing band. Importantly, there is no topological phase in this model. We now consider the same quench protocol as performed for BHZ model discussed in the main text. For a trivial insulator ($|\kappa^{\prime}|\gg 2$), we find that $\Delta E_{j}$ for all the bulk states shows non-universal fluctuation about zero (see Fig. 9(c)). Similar to the trivial phase of BHZ model, here also $\Delta\sigma_{\rm{bulk}}^{f}$ shows non-universal dependence on filling $f$ and $\Delta\sigma_{\rm{bulk}}^{f}<0$ at half-filling (see Fig. 9(d)).
@@
Therefore, the correction to the interfacial line energy $\Delta\sigma_{\rm{bulk}}$ in trivial phase at half-filling (i.e. ground state) can be either positive or negative which is decided by the random contributions from individual single-particle energy-levels, rendering the behavior of the interfacial line energy in trivial situation non-universal.
Appendix B Role of Symmetries and Topological Interface Energy
Here we demonstrate the role of breaking spin-fermion flip symmetry and particle-hole symmetry in affecting the interfacial line energy in topological phases.
@@ Appendix B Role of Symmetries and Topological Interface Energy
B.1 Role of breaking spin-fermion flip symmetry (class D model)
@@
Figure 10: Effect of hopping between next nearest neighbors: (a) Chern number $C$ as a function of mass parameter $M$, (b) change in energy density $\Delta\epsilon$ as a function of $\kappa$, (c) additional interfacial line energy $\Delta\sigma$ as a function of $\kappa$, (d) enhancement ratio $\gamma=n_{c}^{\rm{Q}}/n_{c}^{\rm{cl}}$ of critical cluster size as a function of $\kappa$ when the hopping between next nearest neighboring unit cells is absent ($t_{2}=0.0$) and present ($t_{2}=1.0$). In all plots, $J=0.2$, $h=1.0$, $T=0$, $L^{2}=4096$.
@@
In order to break the spin-fermion flip symmetry (while retaining particle-hole symmetry), we introduce the next-nearest neighbor hopping $t_{2}$ such that the complete spin-fermion hopping Hamiltonian becomes
$H$ $=-\kappa\sum_{i}s_{i}{\bf{\Psi}}_{i}^{\dagger}\sigma_{z}~{\bf{\Psi}}_{i}-\sum_{\langle ij\rangle}({\bf{\Psi}}_{i}^{\dagger}{\large{\eta}}_{ij}~{\bf{\Psi}}_{j}+{\rm{h.c.}})$
$-t_{2}\sum_{\langle\langle ij\rangle\rangle}({\bf{\Psi}}_{i}^{\dagger}{\large{\eta}}^{\prime\prime}_{ij}~{\bf{\Psi}}_{j}+{\rm{h.c.}})-J\sum_{\langle ij\rangle}s_{i}s_{j}-h\sum_{i}s_{i},$ (29)
@@
where $t_{2}$ is the strength of the hopping between the next nearest neighboring unit cells and $\eta^{\prime\prime}_{ij}=\frac{1}{2}(\sigma_{z}+i\cos(\theta_{ij})\sigma_{x}+i\sin(\theta_{ij})\sigma_{y})$, the angle $\theta_{ij}$ between the $x$-axis and the bond connecting the next nearest neighboring unit cells can assume the values $\theta_{ij}=\pi/4$ and $\theta_{ij}=3\pi/4$. For uniform $s_{i}$ throughout the system, we assume $M\equiv\kappa s_{i}$ and the Hamiltonian in momentum space for $k_{x},k_{y}\in(-\pi,\pi]$ is written as
@@
$\sigma_{x}H^{*}_{k_{x},k_{y}}\sigma_{x}=-H_{-k_{x},-k_{y}}.$ (31)
For uniform $s_{i}$, the system can host topological phases having Chern numbers $C=\pm 2,\pm 1$ depending on the value of mass parameter $M$ when $t_{2}\neq 0$ (see Fig. 10(a)). Interestingly, the ground state energy of the fermionic system, which is given by
$E_{g}(M)=\sum_{k_{x},k_{y}}E_{k_{x},k_{y}}$ $=\frac{L^{2}}{(2\pi)^{2}}\int_{-\pi}^{\pi}dk_{x}\int_{-\pi}^{\pi}dk_{y}E_{k_{x},k_{y}}$
$\text{where~}E_{k_{x},k_{y}}$ $=-\sqrt{d_{1}^{2}+d_{2}^{2}+d_{3}^{2}}~,$ (32)
is not symmetric under $M\to-M$, i.e., $E_{g}(M)\neq E_{g}(-M)$ when $t_{2}\neq 0$. Thus, the sudden quench from $h<0$ to $h>0$ causes an additional correction to bulk energy density $\Delta\epsilon$ along with the correction to interfacial line energy $\Delta\sigma$ when $t_{2}\neq 0$, such that the change in the energy in the fermionic system for a formation of a square-shaped droplet of size $n$ is
$\Delta E=(\Delta\epsilon)n+4\sqrt{n}(\Delta\sigma).$ (33)
We show the variation of $\Delta\epsilon$ and $\Delta\sigma$ in Fig. 10(b,c), where we observe that unlike the situation with $t_{2}=0$, $\Delta\epsilon\neq 0$ for $\kappa\neq 0$ when $t_{2}\neq 0$. Furthermore, in the topological situation $0<|\kappa|<2$, $\Delta\sigma$ for $t_{2}\neq 0$ is greater than that for $t_{2}=0$. This can be understood as follows: when $t_{2}\neq 0$ and $0<\kappa<2$, a droplet of $s_{i}=1$ corresponds to Chern number $C=2$, while its background with $s_{i}=-1$ has C=1C=-1. Thus, the difference in Chern numbers ($|\Delta C|=3$) between the droplet and its background for $t_{2}\neq 0$ is greater than $|\Delta C|=2$ for $t_{2}=0$. An increased $|\Delta C|$ in topological situation causes greater $\Delta\sigma$ since the number of edge modes increases, increasing the effective $\Delta{\mathcal{E}}_{\text{edge}}$. This results in a higher $\Delta\sigma$ for $t_{2}\neq 0$ compared to that for $t_{2}=0$. For $2<\kappa<0$, $\Delta\sigma$ shows similar behavior where a droplet of $s_{i}=1$ has C=1C=-1 and its background with $s_{i}=-1$ has $C=2$. In presence of non-zero $\Delta\epsilon$ and $\Delta\sigma$, the critical cluster size for $t_{2}\neq 0$ modifies to
$n_{c}^{\rm{Q}}=\frac{\pi(\sigma_{\rm{cl}}+\Delta\sigma)^{2}}{(2|h|-\Delta\epsilon)^{2}},$ (34)
@@
$\gamma=\frac{n_{c}^{\rm{Q}}}{n_{c}^{\rm{cl}}}=\frac{(1+(\Delta\sigma/\sigma_{\rm{cl}}))^{2}}{(1-(\Delta\epsilon/(2|h|)))^{2}}.$ (35)
We find that in topological situation, $\gamma$ can be greater than $1$ for a suitable choice of $J$ and $h$, indicating an enhancement of critical cluster size (see Fig. 10(d)).
Figure 11: Effect of particle-hole symmetry breaking hopping term: (a) Chern number $C$ as a function of mass parameter $M$, (b) change in energy density $\Delta\epsilon$ as a function of $\kappa$, (c) additional interfacial line energy $\Delta\sigma$ as a function of $\kappa$, (d) enhancement ratio $\gamma=n_{c}^{\rm{Q}}/n_{c}^{\rm{cl}}$ of critical cluster size as a function of $\kappa$ when particle-hole symmetry breaking hopping term is absent ($\lambda=0.0$) and present ($\lambda=0.6$). Dotted lines denote the critical points $\kappa=0,\pm 2$ when $\lambda=0$ and $\kappa=0,\pm 1.6$ when $\lambda=0.6$. In all plots, $J=0.2$, $T=0$, $L^{2}=4096$.
Figure 12: Interfacial line energy and critical cluster-size at finite temperature: (a) Quantum correction to the interfacial line energy $\Delta\sigma$ as a function of $\kappa$ at temperatures $T=0$ and $T=1/\beta=0.33$, (b) $\sigma_{\rm{cl}}$ of classical Ising model ($\kappa=0$) as a function of $T$, (c) $\gamma=n_{c}^{\rm{Q}}/n_{c}^{\rm{cl}}=(1+(\Delta\sigma/\sigma_{\rm{cl}}))^{2}$ as a function of $\kappa$ at temperatures $T=0$ and $T=1/\beta=0.33$, where $n_{c}^{\rm{cl}}$ is the critical cluster-size for classical Ising model ($\kappa=0$). For all the plots, number of unit cells is $L^{2}=4096$. The parameters chosen are: $J=0.2$ (where $T_{c}\approx 0.45$) and $h=0.02$. Here, $\sigma_{\rm{cl}}\sim 2J$ at $T=0$. At finite $T$ in (b), $n_{c}^{\rm{cl}}$ is evaluated using Monte Carlo simulation for classical Ising model ($\kappa=0$) via Eq. (39) and $\sigma_{\rm{cl}}$ is then obtained from $n_{c}^{\rm{cl}}$ using Eq. (40). In (c), $n_{c}^{\rm{cl}}=145$ for $T=1/\beta=0.33$.
B.2 Role of breaking particle-hole symmetry (class A model)